OpenStudy (anonymous):

Can someone please help me understand how to write equations in standard form using integers?

OpenStudy (anonymous):

\[y=2x+5\]

OpenStudy (anonymous):

so standard form is x+y = some number so here we have to get 5 on one side by itself

OpenStudy (anonymous):

Ok so tahts the part that cnfuses me is getting the number 5 alone on 1 side. The quation is adding so that means you'd have to subtract 2x from each side???

OpenStudy (mathstudent55):

yes

OpenStudy (mathstudent55):

Standard form for a linear equation (the equation that graphs as a straight line) is ax + by = c, where a, b, and c are numbers. In this case, subtract 2x from both sides, and you'll get it.

OpenStudy (anonymous):

Thats the part that really confuses me...I got 5=y-2x.

OpenStudy (mathstudent55):

Start again: y = 2x + 5 The y is already on the left, and the 5 is already on the right. All you need to do, is get the 2x on the left side.

OpenStudy (anonymous):

ok so then itd be 2x-y=5.

OpenStudy (mathstudent55):

To do that, subtract 2x from both sides: y = 2x + 5 -2x -2x ----------------- -2x + y = 5 Answer is -2x + y = 5

OpenStudy (anonymous):

Ok thank you very much!

OpenStudy (anonymous):

that's actually the same thing you got the first time though